Dollar/Euro Forwards. Use the following spot and forward​ bid-ask rates for the U.S.​ dollar/euro

​(US$/euro€​)

from December

​10, 2010, to answer the following​ questions:

a. What is the​ mid-rate for each​ maturity?

b. What is the annual forward premium for all​ maturities?

c. Which maturities have the smallest and largest forward​ premiums?

Period

Bid Rate

Ask Rate

spot

1.32311.3231

1.32321.3232

1 month

1.32301.3230

1.32311.3231

2 months

1.32281.3228

1.32291.3229

3 months

1.32241.3224

1.32271.3227

6 months

1.32151.3215

1.32181.3218

12 months

1.31941.3194

1.31981.3198

24 months

1.31471.3147

1.31761.3176

Days Bid Rate Ask Rate Forward Period Forward USSIE US$/€ Premium Spot 1.3231 1.3232 1 month 30 1.3230 1.3231
Answer #1
Mid rate or mid price is simply the average of bid rate and ask rate
Mid rate= (bid rate + ask rate)/2
Period Bid Rate (a) Ask Rate (b) Mid rate ('c)
c= =(a+b)/2
spot 1.3231 1.3232 1.32315
1 month 1.323 1.3231 1.32305
2 months 1.3228 1.3229 1.32285
3 months 1.3224 1.3227 1.32255
6 months 1.3215 1.3218 1.32165
12 months 1.3194 1.3198 1.3196
24 months 1.3147 1.3176 1.31615
Forward premium (or discount) is the difference spot rate and forward rate
Forward premium = ( Forward - Spot )
Forward premium = ( Forward - Spot ) / (Spot) x (360 / days)
Period Bid Rate (a) Ask Rate (b) Mid rate ('c) Forward premium (d) Days to forward/ maturity Annualized forward premium ('e)
c= =(a+b)/2 =forward rate- spot rate =forward premium*360/forward days
spot 1.3231 1.3232 1.3232              -   0 NA
1 month 1.323 1.3231 1.3231    (0.0001) 30 -0.1200%
2 months 1.3228 1.3229 1.3229    (0.0003) 60 -0.1800%
3 months 1.3224 1.3227 1.3226    (0.0006) 90 -0.2400%
6 months 1.3215 1.3218 1.3217    (0.0015) 180 -0.3000%
12 months 1.3194 1.3198 1.3196    (0.0036) 360 -0.3550%
24 months 1.3147 1.3176 1.3162    (0.0070) 720 -0.3500%
The 24 month forward rate has the largest premium (discount = 0.007), while the 1 month forward possesses the smallest premium (discount = 0.0001)